OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. x 2 + (y - 3) 2 = 8

OpenStudy (anonymous):

\((x-h)^2 + (y-k)^2 = r^2\) Where (h,k) is the center and r is the radius. :-) Got it?

OpenStudy (anonymous):

Hint: (any constant number) \(\pm\) 0 = 0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am confused by that answer.

OpenStudy (anonymous):

\(x^2 + (y - 3)^2 = 8\) can be written as \((x−0)^2+(y−3)^2=8\) so it looks like this \((x-h)^2 + (y-k)^2 = r^2\) and you can easily figure out the radius and center (re-read what I wrote above) If r^2 = 8, what does r = ? (hint: What "undoes" a "^2"? Square root yes?) I'm not sure I can get any more explicit than that... Try it!

OpenStudy (anonymous):

thanks